OpenStudy (anonymous):

A woman pushes a grocery cart down the aisle at a distance of 85 feet. She exerts a force of 32 pounds downward at an angle of 15 degrees with the horizontal. How much work does the woman do on the cart?

OpenStudy (anonymous):



OpenStudy (anonymous):

Oops should be \(15^\circ\) anyway... the point is you only want to find \(x\) here. Remember SOH CAH TOA.

OpenStudy (anonymous):



OpenStudy (anonymous):

Really blundered on that one. Anyway, you need to get rid of the downward force and only consider the horizontal force.

OpenStudy (anonymous):

You could also remember that: \[ W = \vec{F}\cdot \vec{d} = || \vec{F}||\ ||\vec{d}||\ \cos(\theta) \]